1- You are the operations manager for an airline and you are considering a higher fare...

1-

You are the operations manager for an airline and you are considering a higher fare level for passengers in aisle seats. How many randomly selected air passengers must you​ survey? Assume that you want to be

9090​%

confident that the sample percentage is within

1.51.5

percentage points of the true population percentage. Complete parts​ (a) and​ (b) below.

a. Assume that nothing is known about the percentage of passengers who prefer aisle seats.

nequals=nothing

​(Round up to the nearest​ integer.)

b. Assume that a prior survey suggests that about

3333​%

of air passengers prefer an aisle seat.

nequals=nothing

​(Round up to the nearest​ integer.)

2-

A food safety guideline is that the mercury in fish should be below 1 part per million​ (ppm). Listed below are the amounts of mercury​ (ppm) found in tuna sushi sampled at different stores in a major city. Construct a

9999​%

confidence interval estimate of the mean amount of mercury in the population. Does it appear that there is too much mercury in tuna​ sushi?

0.500.50  

0.740.74  

0.110.11  

0.980.98  

1.321.32  

0.520.52  

0.890.89

  

What is the confidence interval estimate of the population mean

muμ​?

nothing

ppmless than<muμless than<nothing

ppm

​(Round to three decimal places as​ needed.)

Does it appear that there is too much mercury in tuna​ sushi?

A.

​No, because it is not possible that the mean is greater than 1 ppm.​ Also, at least one of the sample values is less than 1​ ppm, so at least some of the fish are safe.

B.

​Yes, because it is possible that the mean is greater than 1 ppm.​ Also, at least one of the sample values exceeds 1​ ppm, so at least some of the fish have too much mercury.

C.

​No, because it is possible that the mean is not greater than 1 ppm.​ Also, at least one of the sample values is less than 1​ ppm, so at least some of the fish are safe.

D.

​Yes, because it is possible that the mean is not greater than 1 ppm.​ Also, at least one of the sample values exceeds 1​ ppm, so at least some of the fish have too much mercury.
Answer #1

Gunition Given Alisema Corlidence interval frr): 904 09 confidence that the sample perontage is eithiolt1-50% Iof papoution oD: (0-31)(0.67) * ( be bus ..^=(0.33) (0,67)*10006.777 In = 2659.120 ô n 12659 CS Scanned with CamScanner

2nd question

answer

Standard Deviation, s: 0.39101211770286 Count, N: 7 Sum, Ex 5.06 Mean, X 0.72285714285714 Variance, s2 0.15289047619048 Steps

(river dola 6.50 ,0.74, 0.1, 0.78,0·23,0:52,0.3 S 0.391 Construct 92. comdistance intemul estimate of the Pop Jetion mpon M.= (0.145 , 1-271) The CI estimate of the population meam M is, 0.175ppm 2M € 19.271) ppm. CS Scannertion B is correct, CamS

NOTE:if you have any doubts please comment below

please rate positive